OpenStudy (unofficialllyy):

A party rental company has chairs and tables for rent. The total cost to rent 6 chairs and 5 tables is $63. The total cost to rent 2 chairs and 3 tables is $35. What is the cost to rent each chair and each table?

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (unofficialllyy):

I need help setting up the equation

OpenStudy (unofficialllyy):

I don't even know where to start.

OpenStudy (anonymous):

To start this is a system of equations problem. Start by writing out the two equations for each separate problem. Show them here and I will continue helping

OpenStudy (unofficialllyy):

6C+5T=63 2C+3T=35

OpenStudy (anonymous):

Nice work! Now take one of the equations and solve for either variable. I suggest solving for chairs. For example, the resulting equation will look like this C=___+____T

OpenStudy (unofficialllyy):

I'm stuck on solving for c

OpenStudy (unofficialllyy):

So if I have 6c+5t=63 would I subtract by 5t on both sides to get 6c=5t+63?

OpenStudy (anonymous):

Yes that is exactly what you do. Then divide by 6 on both sides to get C by itself.

OpenStudy (unofficialllyy):

I know 6c divded by 6 makes c, but I'm not sure how to divide 6 by 5t+63

OpenStudy (anonymous):

Ok so you will end up with fractions but that is ok. So it would be (5/6)T+(63/6)=C if this makes sense

OpenStudy (unofficialllyy):

Yes

OpenStudy (anonymous):

Ok. Now you can plug this back into the equation for C. So we are going to use the equation 2C+3T=35. Because we know what C is we can have an equation with one variable. When you see C in the equation you can plug in the new equation for T. For example, 2(5/6T+63/6)+3T=35. Does this make sense?

OpenStudy (unofficialllyy):

Yes I solved and got 3 for t but what about c?

OpenStudy (anonymous):

Good now to solve for C you go back to the original equation and plug in 3 for T. For example for the equation 2C+3T=35, you would plug in 2C+3(3)=35 and solve for C.

OpenStudy (unofficialllyy):

Oh I get it now! I got 13. Thank you so much!

OpenStudy (anonymous):

No problem! Im glad I could help